OpenStudy (anonymous):

Calculate y′ x^2 cos(y) + sin(2y) = xy

OpenStudy (anonymous):

Try this website: https://mathway.com/

OpenStudy (abmon98):

Use product rule to differeniate (x^2cos(y))'=2xcos(y)-y'sin(y)x^2 (xy)=y+xy' d/dy sin(2y)=cos(2y)*d/dy(2y)=2y'cos(2y) -y'sin(y)x^2+2y'cos(2y)-xy'=y-2xcos(y) y'(-sin(y)x^2+2cos(2y)-x)=y-2xcos(y) y'=y-2xcos(y)/(-sin(y)x^2+2cos(2y)-x)

OpenStudy (anonymous):

Thanx👍🏼